Order VPS Hosting
Order a VPS, Semi- dedicated or Dedicated server in Dallas, London or Australia.

Get Assistance
Ask our support team about your hosting requirements.

Host where the staff takes pride in making customers happy

I would just like to tell all of you how pleased we are with your service. The setup was great, the response to questions from the support staff has been great. We are very happy to be working with RIMU, as it is making our jobs much easier.

- Allan (in a 'thank you' ticket) (#99/334)
Home > Support > Notices > NZ datacenter issues

Related Links

Notice Links:


NZ datacenter issues

PostedSun, 1 Apr 2012 07:05 AM UTC
Sun, 1 Apr 2012 03:05 AM EDT
Last UpdateSun, 1 Apr 2012 23:07 PM UTC (572 weeks ago)
Sun, 1 Apr 2012 19:07 PM EDT
Affected Servershost807.rimuhosting.com , host808.rimuhosting.com , host851.rimuhosting.com , host867.rimuhosting.com , host963.rimuhosting.com
Affected Data CenterAuckland

A number of our servers in this datacenter are not responding. The on-call technician is presently traveling to the datacenter to investigate and resolve this issue. We will update as soon as more information is on hand.

Update 0749 UTC: Some servers are up already. We are working to bring those online smoothly.

Update 0825 UTC: Most servers are up now, there are a few VPSs starting still.

Update 0830 UTC: The cause of this outage was isolated to the Power board for affected servers failing after a server PSU died. That server has been removed from the rack and normal power is restored.  The data center staff report that the PSU blew the original PDU.  A new PDU was found and it blew that also.  After that PSU was disconnected the current PDU 'held'.

Update 0850 UTC: Seeing some io issues on host851, working on that.

Update 0915 UTC: All VPSs on host851 should be responding now. There is still some io on that though. Investigating host807, seeing a raid issue on this server. Having to restart that again to clear that up. Please bear with us.

Update 0930 UTC: still working on host807, having that restarted manually.

Update 1000 UTC: host807 is back up. A drive was found to have failed and was causing an io storm. That has been removed from member arrays. We will need to replace that during scheduled maintenance. VPSs are starting gradually.

Update 1040 UTC: all VPSs and servers should be back up and running now. If you are still seeing issues please raise a ticket with us.


Keep You Updated?

Log in to subscribe to changes to this notice.

Set your contact details for future notifications.